Question:
3 times the number which is x less than two is eight.
how would you write this as an algebraic equation. Thank you.

Answer by stanbon(75887)

3 times the number which is x less than two is eight.
------------------------
Equation:
3(2-x) = 8
-----
6 - 3x = 8
3x = -2
--------
x = -2/3
=====================
Cheers,
Stan H.